OpenStudy (anonymous):

What is the asymptote of the graph of the function f(x)=-6^x+4

OpenStudy (accessdenied):

The asymptote occurs on one end of the exponential function, where you get larger negative exponents. Notice that: \( 6^x \) For a negative number, say x = -a, a>0 \( 6^{-a} = \dfrac{1}{6^a} \) For larger and larger negative values (large positive a), the denominator gets very very big, and numerator is a measly 1. 1/ huge numbers tends towards a tiny value, near 0. So the end behavior of the function is that 6^x --> 0, and y = -6^x + 4 = 0 + 4